A pole that is 2.8 m tall cast a shadow that is 1.01 m long. At the same time a nearby building cast a shadow that is 44.5 m long. How tall is the building? Round your answer to the nearest meter
Answer by Boreal(15235)

this is a proportion
thing: shadow
2.8/1.01=x/44.5 units m
cross-multiply
1.01x=445*2.8=124.6
x=123.36 m or 123 m